OpenStudy (anonymous):

simplify: 5^-1 x^0 all divided by y^-1

Nnesha (nnesha):

\[\huge\rm \frac{ 5^{-1} x^0 }{ y^{-1} }\] anything to the 0 power is just = 1 for example \[\large\rm (jkli22)^0 =1\] \[\large\rm (Nnesha)^0 = 1\] :D (22)^0 =1

OpenStudy (anonymous):

my answer choices are : 1. 1/5y 2. 5y 3. y/5 4.5

OpenStudy (jhannybean):

Remember: \[x^{-1} \implies \frac{1}{x}\]Also \[\frac{1}{x^{-1}} \implies x^1\]

Nnesha (nnesha):

another exponent rule! \[\huge\rm m^{-1} = \frac{ 1 }{ m }\]

OpenStudy (anonymous):

so is it 1 or 2 or 3 or 4 I got 3

Nnesha (nnesha):

you're right!

OpenStudy (jhannybean):

That's correct.

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

np :-) btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{gold}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)